r2 - 23 Mar 2010 - 13:20:13 - KaushikDeYou are here: TWiki >  Admins Web > MinutesDataManageMar23



Minutes of the US ATLAS Data Management meeting, Mar. 23, 2010
  • Previous meetings and background : IntegrationProgram
  • Coordinates: Tuesdays, Noon Central
    • (605) 715-4900, Access code: 735188; Dial *6 to mute/un-mute.


  • Meeting attendees: Shawn, Bob, Armen, John, Saul, Charles, Aaron, Wensheng, Hiro
  • Apologies: Pedro
  • Guests:

Topics for this week

  • Space status at BNL - Armen
    • Everything looks OK
    • GROUPDISK - 60 TB allocated, 10 TB free. OK.
  • Space status at Tier 2's - Armen
    • MWT2 added another 10 TB of headroom on MCDISK. New storage should come by end of week.
    • NET2 - 40 TB on DATADISK and 10 TB on MCDISK free. Need to free some space, since no new storage for at least next couple of months.
    • Hiro cleaned up USERDISK.
  • FTS 2.2 and SS migration status - Hiro
    • Only NET2 not migrated. Planning to do it today.
    • All other sites working fine.
  • Hot issues
    • Some deletions not taking place on T2 MCDISK - Charles following up.
  • AOB

-- KaushikDe - 23 Mar 2010

About This Site

Please note that this site is a content mirror of the BNL US ATLAS TWiki. To edit the content of this page, click the Edit this page button at the top of the page and log in with your US ATLAS computing account name and password.


Powered by TWiki
This site is powered by the TWiki collaboration platformCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback